# Persistent and anti-Persistent Motion in Bounded and Unbounded Space: Resolution of the First-Passage Problem

### v1

@@ Appendix C Defect Technique in Correlated Random Walks
where $v_{\pm}\left(\frac{2\pi\kappa}{N}\right)=\frac{1}{2b}\left[-f-fe^{\frac{4\pi\kappa}{N}}\pm 2e^{\frac{2\pi\kappa}{N}}\sqrt{b^{2}-f^{2}+f^{2}\cos^{2}\left(\frac{2\pi\kappa}{N}\right)}\right]$, and $\lambda_{\pm}$ are defined in Eq. (9) of the main text. Equation (S23) reduces to Eq. (7) of the main text when $\alpha_{1}=\alpha_{2}$.
@@
The image set used, for the semi-infinite propagator, is dependent on whether the reflection is to the left or the right of the initial condition. We assume, without loss of generality that the barrier is placed between sites $n=0$ and $n=1$ meaning $P(1,\{2,1\},t)=P(0,\{1,2\},t)$.
@@
which is written, after summing the components, as
@@
$\frac{1-cz}{r(z)^{|n+n_{0}-1|}}+\frac{(b-f)z}{2r(z)^{|n+n_{0}|}}+\frac{(b-f)z}{2r(z)^{|n+n_{0}-2|}}\bigg{]}.$
In the ($b=0$) ballistic limit, the boundary condition can be simplified to $P(0,1,t+1)=P(1,2,t)$ as there will never be an instance of the walker, entering state $m=1$ and then crossing the boundary at a later time. In fact, the trajectories in state $m=1$, will remain in $m=1$ for all time, meaning the image of this part of the dynamics will deterministically travel towards minus infinity and never interact with the domain of interest. As such, one can write the propagator in this case as
@@
We derive here the propagator in arbitrary dimensions. Employing the method of images for the reflective case [6],
$\widetilde{\bm{P}}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)=\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}$$\bigg{\{}\widetilde{\bm{Q}}_{n_{0_{1}}+2\kappa_{1}N_{1},...,n_{0_{d}}+2\kappa_{d}N_{d}}(n_{1},...,n_{d},z)$ (S27)
$+\widetilde{\bm{Q}}_{-n_{0_{1}}+1+2\kappa_{1}N_{1},...,-n_{0_{d}}+1+2\kappa_{d}N_{d}}(n_{1},...,n_{d},z)\bigg{\}},$
we obtain, from Eq. (3), the reflective propagator
@@
$\frac{1}{2\pi}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}\left(\left[e^{-i(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\bm{\xi}}+e^{-i(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\bm{\xi}}\right]e^{i\bm{\kappa}^{\intercal}\cdot 2\bm{N}\bm{\xi}}\right)\left[\mathbb{I}-z\bm{\lambda}\left(\xi_{1},...,\xi_{d}\right)\right]^{-1}\cdot\bm{U_{m_{0}}}\text{d}^{d}\xi.$
Using Eq. (S20) with $\mu_{d}=\frac{\pi}{N_{d}}$, one obtains
@@
$\frac{1}{2\pi}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{1}=-\infty}^{\infty}\bigg{(}\delta\left(\bm{\xi}-\frac{\bm{\kappa}}{\bm{N}}\pi\right)\left[e^{-i(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\bm{\xi}}+e^{-(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\bm{\xi}}\right]\bigg{)}\left[\mathbb{I}-z\bm{\lambda}\left(\xi_{1},...,\xi_{d}\right)\right]^{-1}\cdot\bm{U_{m_{0}}}\text{d}\xi,$
where $\frac{\bm{k}}{\bm{N}}=\left[\frac{\kappa_{1}}{N_{1}},...,\frac{\kappa_{d}}{N_{d}}\right]$ which is directly solvable via the same procedure as Eq. (S22) leading to
@@
$\frac{1}{(2N)^{d}}\sum_{\kappa_{1}=0}^{2N_{1}-1}...\sum_{\kappa_{d}=0}^{2N_{d}-1}\left(\big{[}e^{-i\pi(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}+e^{-i\pi(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}\big{]}\right)\left[\mathbb{I}-z\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)\right]^{-1}\cdot\bm{U_{m_{0}}},$
which, after some $z$-inverting, utilising that all the functions are periodic to shift the summation and splitting the summation into parts and isolating the $(\kappa_{1},...,\kappa_{d})=(0,...,0)$-th term, allows passage to
$\bm{P}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},t)=\frac{1}{(2N)^{d}}\underset{\kappa_{1}+\kappa_{2}+...\kappa_{d}>0}{\sum_{\kappa_{1}=0}^{N_{1}-1}...\sum_{\kappa_{d}=0}^{N_{d}-1}}\bigg{\{}\left(\big{[}e^{-i\pi(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}+e^{-i\pi(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}\big{]}\right)\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)^{t}\cdot\bm{U_{m_{0}}}$ (S31)
$+\left(\big{[}e^{i\pi(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}+e^{i\pi(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}\big{]}\right)\bm{\lambda}\left(\frac{-\pi\kappa_{1}}{N_{1}},...,\frac{-\pi\kappa_{d}}{N_{d}}\right)^{t}\cdot\bm{U_{m_{0}}}\bigg{\}}+\frac{2}{(2N)^{d}}\bm{\lambda}(0,...,0)^{t}\cdot\bm{U_{m_{0}}}.$
To proceed we note that while $\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)$ and $\bm{\lambda}\left(\frac{-\pi\kappa_{1}}{N_{1}},...,\frac{-\pi\kappa_{d}}{N_{d}}\right)$ govern different state level dynamics, the site level dynamics are equivalent as the change of sign in the matrix exponent merely alters the state that is entered on each movement direction. Since the propagation is symmetric along all axes (ensured by the uniform initial weighting), then the above reasoning allows one to realise that $\sum_{j=1}^{2d}\left[\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)\cdot\bm{U_{m_{0}}}\right]_{j}=\sum_{j=1}^{2d}\left[\bm{\lambda}\left(\frac{-\pi\kappa_{1}}{N_{1}},...,\frac{-\pi\kappa_{d}}{N_{d}}\right)\cdot\bm{U_{m_{0}}}\right]_{j}$. Hence, upon summation over the states to find the site level propagator
$P^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},t)=\sum_{m=1}^{M}\bm{P}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},t)_{m},$ (S32)
one can factor out the matricial terms in the first two lines of Eq. (S31) leading directly to
$\sum_{m=1}^{M}\bm{P}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},t)_{m}=\sum_{m=1}^{M}\bigg{[}$ $\frac{1}{(2N)^{d}}\underset{\kappa_{1}+\kappa_{2}+...\kappa_{d}>0}{\sum_{\kappa_{1}=0}^{N_{1}-1}...\sum_{\kappa_{d}=0}^{N_{d}-1}}\bigg{\{}\bigg{[}\big{[}e^{-i\pi(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}+e^{-i\pi(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}\big{]}$ (S33)
$+$ $\big{[}e^{i\pi(\bm{n}-\bm{n_{0}})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}+e^{i\pi(\bm{n}+\bm{n_{0}}-\bm{1})^{\intercal}\cdot\frac{\bm{\kappa}}{\bm{N}}}]\bigg{]}\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)^{t}\cdot\bm{U_{m_{0}}}\bigg{\}}+$
$\frac{2}{(2N)^{d}}\bm{\lambda}(0,...,0)^{t}\cdot\bm{U_{m_{0}}}\bigg{]}_{m},$
which via some trigonometric manipulation of the scalar terms in Eq. (S33) allows for direct passage to Eq. (14) with $\gamma=r$ of the main text.
The Squeezing Method
We detail here, via a schematic, the method of transforming a periodic domain into one with reflecting boundaries. The method, first introduced in ref. [45] for the standard LRW case can be applied also to the CRW. The essence of the method can be evinced from Fig. S1. As shown, if the walker occupies site 1 of the periodic lattice, to move off the site its only options are sites 2 and 6. Thus, if we combine sites 2 and 6 together, that is we squeeze the lattice, to create a new site (labelled $2+6$ in the far right panel), we have precisely the reflecting boundary condition considered by Chandrasekhar [50].
Figure S1: (Colour Online). In the style of Fig. 1b in ref. [45], a schematic depicting the method to turn a periodic domain into a reflecting domain by squeezing the internal sites, while the boundary points (sites 1 and 4) remain unaffected by the transformation. For visual clarity we omit arrows depicting the state level dynamics, that is the directions from which a walker may enter a site, depicted in Fig. 1 of the main text. As shown in the right-hand panel, it is clear that after the transformation the walker may enter both internal states from either direction.
@@
Derivation of Correlated Hexagonal Propagator
@@
With the initial condition $\bm{U}_{m_{0}}=\delta_{n_{1},n_{0_{1}}}\delta_{n_{2},n_{0_{2}}}[\alpha_{1},...,\alpha_{6}]^{\intercal}$, and following the techniques in Sec. Derivation of the Unbounded Propagator, , it is possible to write the lattice Green’s function as
$\widetilde{\bm{Q}}^{(\mathcal{H})}_{n_{0}}(\bm{n},z)=\frac{1}{4\pi^{2}}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}e^{-i\bm{\xi}(\bm{n}-\bm{n}_{0})}\bm{\lambda}^{(\mathcal{H})}(\xi_{1},\xi_{2})\cdot\bm{U_{m_{0}}}\text{d}\xi_{1}\text{d}\xi_{2}$ (S44)
where
$\bm{\lambda}^{(\mathcal{H})}(\xi_{1},\xi_{2})=$ (S45)
[…]
with the elements representing the different movement options are labeled, in Eq. (S45), via the arrows.
Figure S2: (Colour Online). A schematic representing two of the six permissible movement steps in the correlated hexagonal random walk where the solid arrows represent the movement probabilities if the walker had entered the site via the dotted arrow. The movements shown on the left side case are shown in the $3^{\text{rd}}$ column of Eq. (S45) and the movements shows in the right panel are encoded in the $2^{\text{nd}}$ column of Eq. (S45).
The Hexagonal Periodic Boundary Condition
$\widetilde{\bm{P}}^{(\mathcal{H})}_{\bm{n}_{0}}(n_{1},n_{2},z)=\sum_{\kappa_{1}=-\infty}^{\infty}\sum_{\kappa_{2}=-\infty}^{\infty}\widetilde{\bm{Q}}_{\bm{n}_{0}}^{(\mathcal{H})}(n_{1}+\hat{n}_{1},n_{2}+\hat{n}_{2},z),$ (S46)
where
$\begin{bmatrix}\widehat{n}_{1}\\ \widehat{n}_{2}\end{bmatrix}=\begin{bmatrix}-R\kappa_{1}+(2R+1)\kappa_{2}\\ -(R+1)\kappa_{1}-R\kappa_{2}\end{bmatrix},$ (S47)
and since again, the correlated periodic boundary conditions are identical to their diffusive counterpart, Eq.  (S44) can be solved explicitly following the same procedure as the solution to the LRW with periodic boundary conditions derived in [55]. This leads to Eq. (20) of the main text.
As mentioned in Sec. IV.2 of the main text, the non-orthogonality of the coordinates leads to a ‘shift’ in the position of the walker when traversing the periodic boundary. For a deeper discussion we again refer the interested reader to [55], and provide here a simple explanation of the dynamics seen in Fig. 3 of the main text. We describe here the path taken by the probability pulse that is near North-East of the domain in the ballistic case Fig. 3(a) (other pulses may then be understood similarly due to the rotational symmetry of the system).
As mentioned in Sec IV.2, this shift leads allows the persistent walker to visit every site in the domain. We show this formally by considering the following [42]. First taking the $z$-transfrom of Eq. (20) such that
$\widetilde{\bm{P}}^{(\mathcal{H})}_{\bm{n}_{0}}(\bm{n},z)=\frac{1}{\Omega}\left[\mathbb{I}-z\bm{\lambda}^{(\mathcal{H})}(0,0)\right]^{-1}+\widetilde{\bm{P}}^{(\kappa\neq 0)}_{\bm{n}_{0}}(\bm{n},z),$ (S48)
where $\widetilde{P}^{(\kappa\neq 0)}_{\bm{n}_{0}}(\bm{n},z)$ is the generating function of the second term in Eq. (20). Since $\bm{\lambda}^{(\mathcal{H})}(0,0)=\mathbb{I}$ for $(\kappa_{1},\kappa_{2})=(0,0)$ only, it is clear that at $z=1$, each element $\widetilde{\bm{P}}^{(\kappa\neq 0)}_{\bm{n}_{0}}(\bm{n},z)$ is holomorphic. Therefore, $\widetilde{\bm{P}}^{(\mathcal{H})}_{\bm{n}_{0}}(\bm{n},z)\to\infty$ as $z\to 1$ for any $\bm{n}$ meaning it is certain all sites will be hit, irrespective of the movement parameters.
@@
Survival Arguments
With a single target we have a set $\mathcal{S}$ of targets with cardinality one, namely $\mathdutchcal{S}=\left\{(\bm{s},m_{\bm{s}})\right\}$. Starting from Eq. (33) of Appendix C in the main text, we find
@@
Renewal Arguments
We start with a set of two coupled equations, which relate the probability of being found in the specific target state and the corresponding first passages
$\alpha_{m_{0_{1}}}P_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},t)$ $=\alpha_{m_{0_{1}}}\delta_{m_{0},1}\delta_{t,0}\delta_{\bm{s},\bm{n}_{0}}+\sum_{t^{\prime}=0}^{t}F_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},t^{\prime})P_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},t-t^{\prime}),$ (S54)
[…]
Assuming $\bm{n}\neq\bm{n}_{0}$ and taking the generating function we find
$\alpha_{m_{0_{1}}}\widetilde{P}_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},z)$ $=\widetilde{F}_{\bm{n}_{0},1}(\bm{s},m_{\bm{s}},z)\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z),$ (S55)
[…]
For each equation in Eq. (S55) one may solve for
$\widetilde{F}_{\bm{n}_{0},i}(\bm{s},m_{\bm{s}},z)=\frac{\alpha_{m_{0_{i}}}\widetilde{P}_{\bm{n}_{0},i}(\bm{s},m_{\bm{s}},z)}{\widetilde{P}_{\bm{s},m_{\bm{s}}}(\bm{s},m_{\bm{s}},z)},$ (S56)
independently of the other equations and upon summation $\widetilde{F}_{\bm{n}_{0}}(\bm{s},m_{\bm{s}},z)=\sum_{i=1}^{M}\widetilde{F}_{\bm{n}_{0},i}(\bm{s},m_{\bm{s}},z)$, we find Eq. (24) of the main text. Evidently one also sees here that if $\alpha_{m_{0_{i}}}=1$ then one recovers the known cases of ref. [35] and ref. [34].
@@
$\widetilde{F}_{n_{0}}(\mathdutchcal{S},z)=\widetilde{F}_{\bm{n}_{0}}(\mathdutchcal{S},z)=\sum_{j=1}^{M}\sum_{i=1}^{S}\alpha_{m_{\bm{s}_{i}}}\frac{\det[\mathbb{H}^{(i)}(\bm{n}_{0},m_{0_{j}},z)]}{\det[\mathbb{H}(z)]}.$ (S57)

### v2

@@ Appendix D Alternative Derivation of the Single Target First-Passage Probability
where $v_{\pm}\left(\frac{2\pi\kappa}{N}\right)=\frac{1}{2b}\left[-f-fe^{\frac{4\pi i\kappa}{N}}\pm 2e^{\frac{2\pi i\kappa}{N}}\sqrt{b^{2}-f^{2}+f^{2}\cos^{2}\left(\frac{2\pi\kappa}{N}\right)}\right]$, and $\lambda_{\pm}$ are defined in Eq. (9) of the main text. Equation (S23) reduces to Eq. (7) of the main text when $\alpha_{1}=\alpha_{2}$.
@@
The image set used, for the semi-infinite propagator, is dependent on whether the reflection is to the left or the right of the initial condition. We assume, without loss of generality, that the barrier is placed between sites $n=1$ and $n=0$ (with $n_{0}\geq 1$) meaning $P(1,\{2,1\},t)=P(0,\{1,2\},t)$.
@@
which is written, after summing the components from Eq. (S17), as
@@
$\frac{1-cz}{r(z)^{|n+n_{0}-1|}}+\frac{(b-f)z}{2r(z)^{|n+n_{0}|}}+\frac{(b-f)z}{2r(z)^{|n+n_{0}-1|}}\bigg{]}.$
In the ($b=0$) ballistic limit, the boundary condition can be simplified to $P(0,1,t)=P(1,2,t)$ as there will never be an instance of the walker, entering state $m=1$ and then crossing the boundary at a later time. In fact, the trajectories in state $m=1$, will remain in $m=1$ for all time, meaning the image of this part of the dynamics will deterministically travel towards minus infinity and never interact with the domain of interest. As such, one can write the propagator in this case as
@@
We derive here the propagator in arbitrary dimensions. The method of images required to satisfy the ‘bouncing’ reflective boundary condition along the $j^{\text{th}}$ dimension may be written as [6],
$\widetilde{\bm{P}}^{(r)}_{n_{0_{1}},...,n_{0_{j}},n_{0_{d}}}(n_{1},...,n_{j},...,n_{d},z)=\sum_{\kappa_{j}=-\infty}^{\infty}$$\bigg{\{}\widetilde{\bm{Q}}_{n_{0_{1}},...,n_{0_{j}}+2\kappa_{j}N_{j},...,n_{0_{d}}}(n_{1},...,n_{j},...,n_{d},z)$ (S27)
$+\widetilde{\bm{Q}}_{n_{0_{1}},...,-n_{0_{j}}+1+2\kappa_{j}N_{j},...,n_{0_{d}}}(n_{1},...,n_{j},...,n_{d},z)\bigg{\}},$
where we have related the position of the walker in any dimension to its image set along that axis independently of any other dimension, since the nearest-neighbour transition probabilities dictate the walker may never interact with two boundaries simultaneously.
Writing the unbounded propagator, Eq. (S28) (also given in Eq. (3) of the main text), as
$\widetilde{\bm{Q}}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)=\frac{1}{(2\pi)^{d}}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\left[\prod_{j=1}^{d}e^{-i\xi_{j}\left(n_{j}-n_{0_{j}}\right)}\right]\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},...,\xi_{d})\right]^{-1}\cdot\bm{U_{m_{0}}}\text{d}^{d}\xi,$ (S28)
and applying the method of images along each dimension independently, we find the reflective propagator as
@@
$\frac{1}{(2\pi)^{d}}\int_{-\pi}^{\pi}...\int_{-\pi}^{\pi}\sum_{\kappa_{1}=-\infty}^{\infty}...\sum_{\kappa_{d}=-\infty}^{\infty}\left[\prod_{j=1}^{d}e^{-i(n_{j}-n_{0_{j}})\xi_{j}}+e^{-i(n_{j}+n_{0_{j}}-1)\xi_{j}}\right]e^{i\bm{\kappa}^{\intercal}\cdot 2\bm{N}\bm{\xi}}\left[\mathbb{I}-z\bm{\lambda}\left(\xi_{1},...,\xi_{d}\right)\right]^{-1}\cdot\bm{U_{m_{0}}}\text{d}^{d}\xi,$
where we have swapped the order of integration and summation as is permissible in theory of generalised functions [46].
Shifting the region of integration from $V=[-\pi,\pi]^{d}$ to $[-\pi-\varepsilon_{j},\pi-\varepsilon_{j}]$ along each dimension and applying Eq. (S20) where now $\mu^{d}=\prod_{j=1}^{d}\frac{\pi}{N_{j}}$ and $\bm{\mu}=\left[\pi/N_{1},...,\pi/N_{d}\right]$, one obtains
@@
$\frac{1}{(2N)^{d}}\int_{-\pi-\varepsilon_{1}}^{\pi-\varepsilon_{1}}\!...\!\int_{-\pi-\varepsilon_{d}}^{\pi-\varepsilon_{d}}\!\sum_{\kappa_{1}=-\infty}^{\infty}\!...\!\sum_{\kappa_{1}=-\infty}^{\infty}\!\delta\left(\bm{\xi}-\bm{\kappa}^{\intercal}\cdot\bm{\mu}\right)\!\left[\prod_{j=1}^{d}e^{-i(n_{j}-n_{0_{j}})\xi_{j}}\!+\!e^{-i(n_{j}+n_{0_{j}}-1)\xi_{j}}\right]\!\left[\mathbb{I}\!-\!z\bm{\lambda}\left(\xi_{1},...,\xi_{d}\right)\right]^{-1}\!\cdot\bm{U_{m_{0}}}\!\text{d}\xi.$
We may now solve the integral via the same procedure as Eq. (S22), that is performing a summation over the singularities of Eq. (S30), leading to
$\widetilde{\bm{P}}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)=\frac{1}{N^{d}}\left[\mathbb{I}-z\bm{\lambda}\left(0,...,0\right)\right]^{-1}\cdot\bm{U_{m_{0}}}+$ (S31)
$\frac{1}{(2N)^{d}}\mathop{\sum_{\kappa_{1}=-(N_{1}-1)}^{N_{1}-1}...\sum_{\kappa_{d}=-(N_{d}-1)}^{N_{d}-1}}_{\kappa_{1}+...+\kappa_{d}\neq 0}\left[\prod_{j=1}^{d}e^{-\frac{i\pi\kappa_{j}(n_{j}-n_{0_{j}})}{N_{j}}}+e^{-\frac{i\pi\kappa_{j}(n_{j}+n_{0_{j}}-1)}{N_{j}}}\right]\left[\mathbb{I}-z\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)\right]^{-1}\cdot\bm{U_{m_{0}}},$
where we have omitted any terms containing the dummy variable $\kappa_{j}=-N_{j}$ as there is clearly no contribution to the summation and have also isolated the $(\kappa_{1},...,\kappa_{d})=(0,...,0)$ term.
Applying Euler’s formula to each term in the product and noticing that the $\sin$ terms give no contribution allows us to write Eq. (S32) as
$\widetilde{\bm{P}}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)=\frac{1}{N^{d}}\left[\mathbb{I}-z\bm{\lambda}\left(0,...,0\right)\right]^{-1}\cdot\bm{U_{m_{0}}}+$ (S32)
$\frac{1}{(2N)^{d}}\!\mathop{\sum_{\kappa_{1}=-(N_{1}-1)}^{N_{1}-1}\!...\!\sum_{\kappa_{d}=-(N_{d}-1)}^{N_{d}-1}}_{\kappa_{1}+...+\kappa_{d}\neq 0}\left[\prod_{j=1}^{d}\cos\left(\frac{\pi\kappa_{j}(n_{j}-n_{0_{j}})}{N_{j}}\right)\!+\!\cos\left(\frac{\pi\kappa_{j}(n_{j}+n_{0_{j}}-1)}{N_{j}}\right)\right]$
×$\quad\quad\quad\quad\times\left[\mathbb{I}-z\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)\right]^{-1}\cdot\bm{U_{m_{0}}}.$
One may simplify this expression by noticing that the periodic nature of the function ensures that for each summation the part in the range $\kappa_{j}\in[-N_{j}+1,-1]$ gives the same contribution as the part in the range $\kappa_{j}\in[1,N_{j}-1]$, and sequentially working from the inside summation outwards, we can write
$\widetilde{\bm{P}}^{(r)}_{n_{0_{1}},...,n_{0_{d}}}(n_{1},...,n_{d},z)=\frac{1}{N^{d}}\left[\mathbb{I}-z\bm{\lambda}\left(0,...,0\right)\right]^{-1}\cdot\bm{U_{m_{0}}}+$ (S33)
$\frac{1}{N^{d}}\!\mathop{\sum_{\kappa_{1}=0}^{N_{1}-1}\!...\!\sum_{\kappa_{d}=0}^{N_{d}-1}}_{\kappa_{1}+...+\kappa_{d}\neq 0}\left[\prod_{j=1}^{d}\cos\left(\frac{\pi\kappa_{j}(n_{j}-n_{0_{j}})}{N_{j}}\right)\!+\!\cos\left(\frac{\pi\kappa_{j}(n_{j}+n_{0_{j}}-1)}{N_{j}}\right)\right]\left[\mathbb{I}-z\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)\right]^{-1}\cdot\bm{U_{m_{0}}}.$
Finally, some using the well-known trigonomentric identity $2\cos(a)\cos(b)=\cos(a+b)+\cos(a-b)$ on each term in the product, taking the inverse $z$-transform and recombining the $(\kappa_{1},...,\kappa_{d})=(0,...,0)$ term, we find
@@
$\frac{1}{N^{d}}\!\sum_{\kappa_{1}=0}^{N_{1}-1}\!...\!\sum_{\kappa_{d}=0}^{N_{d}-1}\left[\prod_{j=1}^{d}\beta_{\kappa_{j}}\cos\left(\frac{\pi\kappa(n_{j}-\frac{1}{2})}{N}\right)\cos\left(\frac{\pi\kappa(n_{0_{j}}-\frac{1}{2})}{N}\right)\right]\left[\mathbb{I}-z\bm{\lambda}\left(\frac{\pi\kappa_{1}}{N_{1}},...,\frac{\pi\kappa_{d}}{N_{d}}\right)\right]^{-1}\cdot\bm{U_{m_{0}}},$
where $\beta_{\kappa_{j}}=2$ if $\kappa_{j}=0$ otherwise $\beta_{\kappa_{j}}=1$, as reported in Eq. (14) of the main text.
@@
The Governing Equations for the Correlated Hexagonal Propagator
@@
With a single target we have a set $\mathcal{S}$ of targets with cardinality one, namely $\mathdutchcal{S}=\left\{(\bm{s},m_{\bm{s}})\right\}$. Starting from Eq. (24) of Appendix D in the main text, we find
@@
$\widetilde{F}_{\bm{n}_{0}}(\mathdutchcal{S},z)=\sum_{j=1}^{M}\sum_{i=1}^{S}\alpha_{m_{\bm{s}_{i}}}\frac{\det[\mathbb{H}^{(i)}(\bm{n}_{0},m_{0_{j}},z)]}{\det[\mathbb{H}(z)]}.$ (S50)